# RDA-INR: Riemannian Diffeomorphic Autoencoding via Implicit Neural Representations

### v2

@@ 5.2 INRs for LDDMM statistical latent modeling
To illustrate the benefit of INRs for topology-preserving latent modeling, we use the liver dataset and compare our method to DAE [16]. We only compare to this method as this is the only neural network-based LDDMM latent model that is designed to work for meshes.
@@
Figure 10: From left to right: the original mesh data, the downsampled version, and the downsampled smooth version.
@@
First, we evaluate the learned templates trained on the downsampled data. These learned templates can be found in Figure 11. The template learned by our model resembles the template that is learned using the original data, which is depicted in Figure 4(a). Hence, even with less detailed meshes we can learn a high-resolution template. Furthermore, while the DAE template found using the non-smooth data looks non-smooth, the template found with the smooth data looks quite smooth. The latter looks similar to the liver template found with the RDA-INR model. However, the template found by DAE is less smooth than the template found by our model. The main reason for this is the smooth structure of the INR parameterizing the Fréchet mean. In DAE the Fréchet mean is calculated by Deformetrica [87]. For Deformetrica we take a single sample from the dataset as initialization of the Fréchet mean. Hence, we inherit the structure of this initialization, possibly causing a problem with the template.
@@
Figure 11: The learned templates of RDA-INR and DAE on the downsampled datasets in Figure 10.
@@
Figure 12: Reconstructions of a specific test sample using models trained on different variants of the original liver data.
Table 6: Train and test evaluation of RDA-INR and DAE. The table presents the average values and medians (between brackets)of the Chamfer Distance (CD) and Earth Mover distance (EM). The values are obtained by comparing the original data to the reconstructions obtained using the dataset the model is trained on. These values are calculated for the shapes in the train and test splits of the liver dataset. The Chamfer Distance values are of the order $10^{-4}$. The best (smallest) values for the downsampled data are in bold.
@@ 6 Conclusion and future work
Several works recently developed neural network models regarding diffeomorphic registration, such as models for pairwise and groupwise registration, atlas building, and data variability modeling. We presented an overview of the current literature on neural networks for diffeomorphic registration and categorized the works depending on several features. This categorization highlights several research gaps. For instance, the LDDMM-based statistical latent modeling approaches are not resolution independent. Moreover, resolution-independent neural network algorithms for joint shape encoding and groupwise registration do not use the Riemannian geometry of shape space, which is a crucial component of LDDMM PGA. Consequently, these latent space models do not provide insights about shape Fréchet means, geodesics, Riemannian distances between shapes, and data variance.
We have addressed the aforementioned two limitations to highlight the importance of resolution independence and LDDMM physical consistency. Specifically, we presented an INR-based latent model inspired by LDDMM PGA. This resolution-independent model is a deformable template model that solves the Fréchet mean finding problem in LDDMM-based PGA. The main ingredient in this model is the Riemannian regularization on the neural network that deforms the template. Although our model also applies to images, we mainly focus on the case of point cloud and mesh data.
First, we discussed the importance of LDDMM physical consistency by comparing our model to the model with a non-Riemannian pointwise regularization on the deformation. We show that the Riemannian regularization is necessary for the model to perform a proper mean-variance analysis. In other words, our model allows calculating a Fréchet mean of the data, obtaining geodesics and approximating the distance between the template and another object, and estimating the data variance. Furthermore, we demonstrate that the Riemannian regularization can improve the reconstruction of an object and the stability of the reconstruction procedure. In other words, the Riemannian regularization induces a prior that enables us to find more stable factors of variation. Finally, we assess the importance of resolution-independence by comparing our model to another neural network model inspired by LDDMM PGA. We demonstrated improved template learning and improved reconstructions.
In summary, we show how shape and image analysis, Riemannian geometry, and deep learning can be connected. This connection paves the way to more research into how these different disciplines can reinforce each other.
@@ 6.1 Future work
An example of such research for our deformable template model is related to our model’s template deformations. Our learned template deformations constitute geodesics between the template and the reconstructed objects. This makes it possible to obtain template deformations that fit a prior and to reconstruct objects using the end point of such a deformation. However, for some objects in the geodesic, there might not exist a latent code and corresponding reconstruction. As these objects fit a modeling prior, we would like them to correspond to a latent code. To achieve this objective, we might employ the velocity vector field parameterization from Lüdke et al. [19]. In addition, even though our model can estimate the distance between the Fréchet mean and the training data via the aforementioned geodesics, we do not use this for generative modeling. In other words, we do not use the distance calculation for defining a Riemannian distribution on the latent space.
Moreover, we focus on the latent modeling capabilities of our model and not the registration capabilities. Hence, it might be interesting to compare our model to neural network models for pairwise and groupwise registration. As these methods often use the more expressive resolution-dependent U-NET architectures, we hypothesize that these methods would still outperform our latent model. We believe it would be interesting to research resolution-independent neural operator variants of our model that more resemble the U-NET architectures. Finally, our model is not only applicable to shapes but also to images. In future work, we would like to assess our model on image data as well.
Conflict of interest statement
The authors declare that the research was conducted in the absence of any commercial or financial relationships that could be construed as a potential conflict of interest.
Data availability statement
The data that support the findings of this study are currently only available upon reasonable request from the authors. In the appendices, we discuss the datasets used in the manuscript. Instructions on how to obtain the datasets will be provided in a GitHub repository that will be made publicly available in the future. This GitHub repository also contains the code used to perform the numerical experiments.
@@ Appendix A Riemannian geometry
Definition 2 (Tangent space).
Assume we have a smooth curve $\gamma:\mathbb{R}\rightarrow M$ with $\gamma(0)=p$. Define the directional derivative operator at $p$ along $\gamma$ as
@@
Definition 3 (Riemannian manifold).
@@
$d_{M}(p,q)=\min_{\gamma}$ $\int_{0}^{1}\sqrt{g_{\gamma(t)}(\dot{\gamma}(t),\dot{\gamma}(t))}\mathrm{d}t$ (A.1)
@@
where $\dot{\gamma}(t)$ is the tangent vector at $\gamma(t)$ generated by $\gamma$. In this case, Equation (A.1) defines a Riemannian distance on $M$ and the minimizer $\gamma^{*}$ is called a (Riemannian) geodesic.
Remark 3.
In case $M=\mathbb{R}^{d}$ and $g_{p}(a,b)=\langle a,b\rangle_{2}$, we have $d_{M}(p,q)=\lVert p-q\rVert_{2}$ and $\gamma^{*}(t)=p+t(q-p)$.
@@
Definition 4 (Fréchet mean).
@@
Definition 5 (Geodesic submanifold).
@@
Appendix B Occupancy data versus point cloud data
@@
In Figures B.1 and B.2, we see the learned templates and an example of a reconstructed training shape, respectively.
@@
Figure B.1: The learned template of the model learned using occupancy functions (left) and the model learned using the point cloud data (right).
@@
Figure B.2: The reconstructions of a particular training shape when using the model learned using occupancy functions (left) and the model learned using the point cloud data (right).
Figure B.1 shows that both templates resemble a square. Figure B.2 demonstrates that the reconstructions with the occupancy data are worse than the reconstructions with the point cloud data. One possible explanation is that the point cloud loss function encourages the shape’s points to lie on the zero level set of the implicit representation. When using occupancy values, we focus more on the domain around the shape and train on uniformly sampled occupancy values to regress the occupancy function. Hence, when working with point clouds, the emphasis is on the shape itself, rather than the surrounding domain, which is the case with occupancy functions. The focus on the shape itself makes it possible to better reconstruct its details. As the point cloud method yields better reconstructions, we use this method for the numerical experiments.
@@ Appendix C Killing energy as standard LDDMM norm
$\lVert\nu\rVert_{V}^{2}=\int_{\Omega}\frac{1}{2}\lVert J\nu+(J\nu)^{T}\rVert_{F}^{2}+\eta\lVert\nu\rVert_{2}^{2}\mathrm{d}x,$ (C.1)
@@
$\int_{\partial\Omega}\langle\nu,(J\nu+(J\nu)^{T})n\rangle_{l_{2}}\mathrm{d}x=0,$ (C.2)
where $n$ is the normal to the boundary. Both assumptions are, for instance, satisfied when $V\subset C_{0}^{2}(\Omega,\mathbb{R}^{d})$ where $C_{0}^{2}(\Omega,\mathbb{R}^{d})$ is the space of twice continuously differentiable vector fields $\nu$ on the open and bounded domain $\Omega\subset\mathbb{R}^{d}$ such that both $\nu$ and its Jacobian $J\nu$ vanish on $\partial\Omega$ and at infinity.
@@
[…]
Proof.
First, define $\nu_{i}$ as the $i$-th component of the vector field $\nu$. As the Frobenius norm comes from an inner product and $\lVert A\rVert_{F}^{2}=\lVert A^{T}\rVert_{F}^{2}$, we obtain
$\frac{1}{2}\lVert J\nu+(J\nu)^{T}\rVert_{F}^{2}$ $=\lVert J\nu\rVert_{F}^{2}+\langle J\nu,(J\nu)^{T}\rangle_{F}$
@@
$\int_{\Omega}\frac{1}{2}\lVert J\nu+(J\nu)^{T}\rVert_{F}^{2}\mathrm{d}x=\int_{\Omega}\sum_{i=1}^{d}\left\langle\nabla\nu_{i},\nabla\nu_{i}+\frac{\partial}{\partial x_{i}}\nu\right\rangle_{l_{2}}\mathrm{d}x=\sum_{i=1}^{d}\int_{\Omega}\left\langle\nabla\nu_{i},\nabla\nu_{i}+\frac{\partial}{\partial x_{i}}\nu\right\rangle_{l_{2}}\mathrm{d}x.$
@@
$\int_{\Omega}\frac{1}{2}\lVert J\nu+(J\nu)^{T}\rVert_{F}^{2}\mathrm{d}x=\sum_{i=1}^{d}\int_{\Omega}\nabla\cdot\left(\nu_{i}\left(\nabla\nu_{i}+\frac{\partial}{\partial x_{i}}\nu\right)\right)-\nu_{i}\left(\Delta\nu_{i}+\nabla\cdot\left(\frac{\partial}{\partial x_{i}}\nu\right)\right)\mathrm{d}x.$
@@
$\int_{\Omega}\frac{1}{2}\lVert J\nu+(J\nu)^{T}\rVert_{F}^{2}\mathrm{d}x=\sum_{i=1}^{d}\left(\int_{\partial\Omega}\left\langle\nu_{i}\left(\nabla\nu_{i}+\frac{\partial}{\partial x_{i}}\nu\right),n\right\rangle_{l_{2}}\mathrm{d}x-\int_{\Omega}\nu_{i}\left(\Delta\nu_{i}+\nabla\cdot\left(\frac{\partial}{\partial x_{i}}\nu\right)\right)\mathrm{d}x\right).$
@@
$\int_{\Omega}\frac{1}{2}\lVert J\nu+(J\nu)^{T}\rVert_{F}^{2}\mathrm{d}x=\int_{\partial\Omega}\left\langle\nu,\left(J\nu+(J\nu)^{T}\right)n\right\rangle_{l^{2}}\mathrm{d}x-\int_{\Omega}\left\langle\nu,\left(\Delta+\nabla\cdot\nabla^{T}\right)\nu\right\rangle_{l^{2}}\mathrm{d}x.$
Finally, using our assumption in Equation (C.2) gives:
$\int_{\Omega}\frac{1}{2}\lVert J\nu+(J\nu)^{T}\rVert_{F}^{2}\mathrm{d}x=-\int_{\Omega}\left\langle\left(\Delta+\nabla\cdot\nabla^{T}\right)\nu,\nu\right\rangle_{l^{2}}\mathrm{d}x.$
Using the above in combination with Equation (C.1), we get:
$\lVert\nu\rVert_{V}^{2}$ $=\int_{\Omega}\frac{1}{2}\lVert J\nu+(J\nu)^{T}\rVert_{F}^{2}+\eta\lVert\nu\rVert_{2}^{2}\mathrm{d}x$

### v3

@@ 5.2 INRs for LDDMM statistical latent modeling
To illustrate the benefit of INRs for topology-preserving latent modeling, we use the liver dataset and compare our method to the diffeomorphic autoencoder (DAE) [14], which works directly on meshes. This method parameterizes a latent space of discrete velocity vector fields on a rectangular grid and obtains a time-dependent discrete velocity vector field via the linear interpolation from the origin to some other latent vector. This time-dependent velocity vector field deforms a learnable template mesh to obtain shape reconstructions. Image representations like signed distance functions for contours or shapes are not always available and non-trivial to obtain from meshes or point clouds. Hence, we can not easily compare our more general framework to LDDMM-based statistical latent modeling methods, like Hinkle et al. [36], that have this requirement.
@@
Figure 10: Liver training data. From left to right: the original mesh data, the downsampled version, and the downsampled smooth version.
@@
First, we evaluate the learned templates trained on the downsampled data. These learned templates can be found in Figure 11. The template learned by our model resembles the template that is learned using the original data, which is depicted in Figure 5a. Hence, even with less detailed meshes we can learn a high-resolution template. Furthermore, while the DAE template found using the non-smooth data looks non-smooth, the template found with the smooth data looks quite smooth. The latter looks similar to the liver template found with the RDA-INR model. However, the template found by DAE is less smooth than the template found by our model. The main reason for this is the smooth structure of the INR parameterizing the Fréchet mean. In DAE the Fréchet mean is calculated by Deformetrica [15]. For Deformetrica we take a single sample from the dataset as initialization of the Fréchet mean. Hence, we inherit the structure of this initialization, possibly causing a problem with the template.
@@
Figure 11: Learned templates. The learned templates of RDA-INR and DAE on the downsampled datasets in Figure 10. RDA-INR is robust against limited data quality (subsampling artifacts, discretization).
@@
Figure 12: Super resolution reconstruction. Test sample reconstruction on different variants of the original liver data. RDA-INR achieves similar reconstructions independent of resolution.
Table 5: Quantitative reconstruction performance. The table presents the average values and medians (between brackets) of the Chamfer Distance (CD) and Earth Mover distance (EM). The values are obtained by comparing the original data to the reconstructions obtained using the dataset the model is trained on. These values are calculated for the shapes in the train and test splits of the liver dataset. The CD values are of the order $10^{-4}$. The best (smallest) values for the downsampled data are in bold.
@@ 6 Conclusion and future work
Several works recently developed neural network models regarding diffeomorphic registration, such as models for pairwise and groupwise registration, atlas building, and data variability modeling. For instance, there exist LDDMM-based statistical latent modeling approaches, which have the limitation that they are not resolution independent. Moreover, resolution-independent neural network algorithms exist for joint shape encoding and groupwise registration. However, they do not use the Riemannian geometry of shape space, which is a crucial component of LDDMM PGA. Consequently, these latent space models do not provide insights about shape Fréchet means, geodesics, Riemannian distances between shapes, and data variance.
We have addressed the aforementioned two limitations to highlight the importance of resolution independence and physical consistency. Specifically, we addressed these issues in the setting of shapes represented by point clouds or meshes. In this setting, we presented a resolution-independent latent model based on INRs and inspired by LDDMM-based PGA. The main ingredient in this deformable template model is the Riemannian regularization on the neural network that deforms the template. Furthermore, to the best of our knowledge, our latent model is the first in the registration literature using image-based (implicit neural) data representations for dealing directly with point cloud or mesh data.
We use the model to provide a study on the benefits of resolution independence and LDDMM physical consistency. First, we discussed the importance of LDDMM physical consistency by comparing our model to the model with a non-Riemannian pointwise regularization on the deformation. We show that the Riemannian regularization is necessary for the model to perform a proper mean-variance analysis. In other words, our model allows calculating a Fréchet mean of the data, obtaining geodesics and approximating the distance between the template and another object, and estimating the data variance. Furthermore, we demonstrate that the Riemannian regularization can improve the reconstruction of an object and the stability of the reconstruction procedure. In other words, the Riemannian regularization induces a prior that enables us to find more stable factors of variation. Finally, we assess the importance of resolution-independence by comparing our model to another neural network model inspired by LDDMM PGA. We demonstrated improved template learning and improved reconstructions.
In summary, we show how shape and image analysis, Riemannian geometry, and deep learning can be connected. This connection paves the way for more research into how these different disciplines can reinforce each other.
@@ 6.1 Future work
Our learned template deformations constitute geodesics between the template and the reconstructed objects. This makes it possible to obtain template deformations that fit a prior and to reconstruct objects using the end point of such a deformation. However, for some objects in the geodesic, there might not exist a latent code and corresponding reconstruction. A future extension of our framework could be to add this missing geodesic consistency to enforce structure on the latent space. This structure increases interpretability as generally changes in latent space have an a-priori unclear effect on the reconstruction. To add the geodesic consistency, we might employ the velocity vector field parameterization from Lüdke et al. [45].
Moreover, a related future impact of the model could be its use for generative modeling. Based on the Riemannian distance, Fréchet means, and geodesics, we can design a Riemannian probability distribution on the latent space. This reveals a part of the latent space’s structure, making it more interpretable and easier to navigate.
Our model is not only applicable to point cloud and mesh data but also to images. In future work, we would like to apply our model to image data towards understanding the importance of data representation in this framework.
Acknowledgments
The authors would like to thank the associate editor and the two anonymous referees for their time and effort to review our manuscript thoroughly and for their suggestions on strengthening the reported work.
@@ Appendix A Riemannian geometry
Definition A.1 (Tangent space).
Assume we have a smooth curve $\gamma:\mathbb{R}\rightarrow M$ on a manifold $M$ with $\gamma(0)=p$. Define the directional derivative operator at $p$ along $\gamma$ as
@@
Definition A.2 (Riemannian manifold).
@@
(17) $d_{M}(p,q)=\min_{\gamma}$ $\int_{0}^{1}\sqrt{g_{\gamma(t)}(\dot{\gamma}(t),\dot{\gamma}(t))}\mathrm{d}t$
@@
where $\dot{\gamma}(t)$ is the tangent vector at $\gamma(t)$ generated by $\gamma$. In this case, Equation (17) defines a Riemannian distance on $M$ and the minimizer $\gamma^{*}$ is called a (Riemannian) geodesic.
Remark A.3.
In case $M=\mathbb{R}^{d}$ and $g_{p}(a,b)=\langle a,b\rangle_{2}$, we have $d_{M}(p,q)=\left\lVert p-q\right\rVert_{2}$ and $\gamma^{*}(t)=p+t(q-p)$.
@@
Definition A.4 (Fréchet mean).
@@
Definition A.5 (Geodesic submanifold).
@@
Appendix B Additional experiments
B.1 Occupancy data versus point cloud data
@@
In Figures 13 and 14, we see the learned templates and an example of a reconstructed training shape, respectively.
@@
Figure 13: Learned templates. The template of the model learned with occupancy functions (left) and the model learned with the point cloud data (right). Both templates resemble the rectangles data.
@@
Figure 14: Training shape reconstruction. The reconstructions of a particular training shape when using the model learned with occupancy functions (left) and the model learned with the point cloud data (right). The model with the point cloud data obtains a higher quality reconstruction.
Figure 13 shows that both templates resemble a square. Figure 14 demonstrates that the reconstructions with the occupancy data are worse than the reconstructions with the point cloud data. One possible explanation is that the point cloud loss function encourages the shape’s points to lie on the zero level set of the implicit representation. When using occupancy values, we focus more on the domain around the shape and train on uniformly sampled occupancy values to regress the occupancy function. The focus on the shape itself makes it possible to better reconstruct its details. As the point cloud method yields better reconstructions, we use this method for the numerical experiments.
B.2 Comparison RDA-INR and NDF
In the main manuscript, we note that our model can be interpreted as a modification of Neural Diffeomorphic Flow (NDF) [63]. For instance, we modify the data attachment such that we can work with point clouds, modify the deformation regularization functional, and modify how we solve the ordinary differential equations. This section studies the effect of all changes on the reconstruction performance by comparing RDA-INR to NDF on the liver dataset.
We run NDF on the liver dataset, which we preprocess according to the NDF paper. We used the same training parameters the authors used except that we put the latent dimension to $32$ and increased the number of epochs to $3000$. For evaluating NDF, we scale the ground truth meshes, ground truth point clouds, and test reconstructions with a scale factor of $0.75$, which we also did on the training data for RDA-INR, as explained in Appendix D.4. We copy the results of the RDA-INR model trained with Riemannian regularization from the main manuscript.
Table 6 presents the reconstruction metrics on the test set of the liver dataset. RDA-INR and NDF show similar performance based on the Chamfer Distance. However, based on the Earth Mover distance, RDA-INR achieves better results. Hence, we conclude that the changes to NDF do not deteriorate its performance or even improve it when considering the Earth Mover distance.
Table 6: Reconstruction performance. RDA-INR and NDF are evaluated by reconstructing the test set of the liver dataset. We calculate the average value and median (between brackets)of the Chamfer Distance (CD) and Earth Mover distance (EM). The CD values are of the order $10^{-4}$. The best (smallest) values are in bold.
Model Metric
(CD) (EM)
NDF $5.61$ $(4.97)$ $0.0499$ $(0.0479)$
RDA-INR $\boldsymbol{5.31}$ $(\boldsymbol{3.68})$ $\boldsymbol{0.0385}$ $(\boldsymbol{0.0340})$
@@ Appendix C Killing energy as standard LDDMM norm
(18) $\left\lVert\nu\right\rVert_{V}^{2}=\int_{\Omega}\frac{1}{2}\left\lVert J\nu+(J\nu)^{T}\right\rVert_{F}^{2}+\eta\left\lVert\nu\right\rVert_{2}^{2}\mathrm{d}x,$
@@
(19) $\int_{\partial\Omega}\langle\nu,(J\nu+(J\nu)^{T})n\rangle_{l_{2}}\mathrm{d}x=0,$
where $n$ is the normal to the boundary. Both assumptions are, for instance, satisfied when $V\subset C_{0}^{2}(\Omega,\mathbb{R}^{d})$ where $C_{0}^{2}(\Omega,\mathbb{R}^{d})$ is the space of twice continuously differentiable vector fields $\nu$ on the open and bounded domain $\Omega\subset\mathbb{R}^{d}$ such that both $\nu$ and its Jacobian $J\nu$ vanish on $\partial\Omega$.
@@
$\left\lVert\nu\right\rVert_{V}^{2}=\int_{\Omega}\langle(\eta\,{\text{id}}-\Delta-\nabla\cdot\nabla^{T})\nu,\nu\rangle_{l_{2}}\mathrm{d}x=\int_{\Omega}\langle\tilde{L}\nu,\nu\rangle_{l_{2}}\mathrm{d}x,$
where $\tilde{L}:=\eta\,{\text{id}}-\Delta-\nabla\cdot\nabla^{T}$.
Proof C.2.
First, define $\nu_{i}$ as the $i$-th component of the vector field $\nu$. As the Frobenius norm comes from an inner product and $\left\lVert A\right\rVert_{F}^{2}=\left\lVert A^{T}\right\rVert_{F}^{2}$, we obtain
$\frac{1}{2}\left\lVert J\nu+(J\nu)^{T}\right\rVert_{F}^{2}$ $=\left\lVert J\nu\right\rVert_{F}^{2}+\langle J\nu,(J\nu)^{T}\rangle_{F}$
@@
$\int_{\Omega}\frac{1}{2}\left\lVert J\nu+(J\nu)^{T}\right\rVert_{F}^{2}\mathrm{d}x=\int_{\Omega}\sum_{i=1}^{d}\left\langle\nabla\nu_{i},\nabla\nu_{i}+\frac{\partial}{\partial x_{i}}\nu\right\rangle_{l_{2}}\mathrm{d}x=\sum_{i=1}^{d}\int_{\Omega}\left\langle\nabla\nu_{i},\nabla\nu_{i}+\frac{\partial}{\partial x_{i}}\nu\right\rangle_{l_{2}}\mathrm{d}x.$
@@
$\int_{\Omega}\frac{1}{2}\left\lVert J\nu+(J\nu)^{T}\right\rVert_{F}^{2}\mathrm{d}x=\sum_{i=1}^{d}\int_{\Omega}\nabla\cdot\left(\nu_{i}\left(\nabla\nu_{i}+\frac{\partial}{\partial x_{i}}\nu\right)\right)-\nu_{i}\left(\Delta\nu_{i}+\nabla\cdot\left(\frac{\partial}{\partial x_{i}}\nu\right)\right)\mathrm{d}x.$
@@
$\int_{\Omega}\frac{1}{2}\left\lVert J\nu+(J\nu)^{T}\right\rVert_{F}^{2}\mathrm{d}x=\sum_{i=1}^{d}$ $\left(\int_{\partial\Omega}\left\langle\nu_{i}\left(\nabla\nu_{i}+\frac{\partial}{\partial x_{i}}\nu\right),n\right\rangle_{l_{2}}\mathrm{d}x-\right.$
$\,\,\,\,\left.\int_{\Omega}\nu_{i}\left(\Delta\nu_{i}+\nabla\cdot\left(\frac{\partial}{\partial x_{i}}\nu\right)\right)\mathrm{d}x\right).$
@@
$\int_{\Omega}\frac{1}{2}\left\lVert J\nu+(J\nu)^{T}\right\rVert_{F}^{2}\mathrm{d}x=\int_{\partial\Omega}\left\langle\nu,\left(J\nu+(J\nu)^{T}\right)n\right\rangle_{l^{2}}\mathrm{d}x-\int_{\Omega}\left\langle\nu,\left(\Delta+\nabla\cdot\nabla^{T}\right)\nu\right\rangle_{l^{2}}\mathrm{d}x.$
Finally, using our assumption in Equation (19) gives:
$\int_{\Omega}\frac{1}{2}\left\lVert J\nu+(J\nu)^{T}\right\rVert_{F}^{2}\mathrm{d}x=-\int_{\Omega}\left\langle\left(\Delta+\nabla\cdot\nabla^{T}\right)\nu,\nu\right\rangle_{l^{2}}\mathrm{d}x.$
Using the above in combination with Equation (18), we get:
$\left\lVert\nu\right\rVert_{V}^{2}$ $=\int_{\Omega}\frac{1}{2}\left\lVert J\nu+(J\nu)^{T}\right\rVert_{F}^{2}+\eta\left\lVert\nu\right\rVert_{2}^{2}\mathrm{d}x$
@@
The theorem shows that our norm, together with the regularization constant $\sigma$, corresponds to the LDDMM norm with $L=(\text{id}-\alpha\Delta-b\nabla\cdot\nabla^{T})^{c}$ for some $\alpha=b>0$ and $c=1$. Interestingly, in the few instances that LDDMM literature uses the $b>0$ case, only Yang et al. [78] use $c=1$, while others typically use $c=2$. While Yang et al. [78] do not provide modeling for the $c=1$ case, in the $c=2$ case, the term with $b>0$ stems from fluid mechanics modeling, with the identity mapping introduced to ensure that the operator is invertible.
Our approach differs in motivation by incorporating the Killing energy to induce a rigid body motion prior, which contrasts with the fluid mechanics perspective seen in other works. Therefore, our method, which results in the $c=1$ case, is novel. Additionally, the $\nabla\cdot\nabla^{T}$ term in our regularizer can be viewed as a rigidity extension of the more common LDDMM operator $L=(\text{id}-\alpha\Delta)^{c}$ when using $c=1$.